2. Given the following data representing the goods market in an open economy Marginal propensity to consume Autonomous consumption Direct tax rate Autonomous Taxation Transfers Gov spending 0.5 300 0.1 100 200 Investment Mareinal propensity to import Autonomous imports Exports 10000 2000 0.2 50 6000 uning the Keymnesan cross model compute a) The equilibrium level of the aggregate output b) The value of public savings corresponding to the equilbrium level of the output; say if the country is running a budge deficit or surplus The value of net exports corresponding to the equilibrium level of the output; say if the country is running a trade deficit or surplus c) d) The value of the agcregate output for which the system would run a trade balance; e) The variation in the equlibrium level of the aggregate output due to a variation in the government spending equal to AGs 1000 The new eqǐbnum level of the aggregate output if the marginal propensity to import decreases to 0.1.
Answer #1

C = 300 + 0.5(Y - 0.1Y- 100 + 200 )

G =. 10,000

I = 2000

IMPORTS = 50 + 0.2Y

EXPORTS = 6000

a) Y = C + I + G + (X- M)

Y = 300 + 0.5(Y - 0.1Y- 100 + 200 ) + 2000 + 10,000 + 6000 - 50 -  0.2Y

Y= 18,250 + 0.5Y - 0.05Y - 50 + 100 - 0.2Y

Y= 18,300 - .25Y

.75Y = 18,300

Y = 24,400

B ) G = 10,000

transfers = 200

TAXES = 0.1Y + 100 = 0.1(24,400 ) + 100 = 2440 + 100 = 2540

since receipts (taxes )< expenditure ( government spending + transfers ) , government is running. a budget deficit

deficit = G + transfers - taxes

= 10,000 + 200 - 2540

= 7660

c) net exports = exports - imports

imports = 50 + 0.2Y = 50 + 0.2*24,400 = 50 + 4880 = 4930

exports = 6000

so net exports = 6000 - 4930

= 1070

since net exports are positive , there is a trade surplus

d) trade balance would occur when exports = imports

= 6000 = 50 + 0.2Y

5950 = 0.2Y

Y = 29,750

trade balance would occur when aggregate output = 29,750

(answered first 4 parts as per policy )